# C-MORL: Multi-Objective Reinforcement Learning through Efficient Discovery of Pareto Front

### v1

@@ Appendix A Policy Selection and C-MORL Algorithm
In this Section, we provide a brief proof for the convergence of iterates majorly inspired by [Tessler et al., 2018]. For detailed proof, we refer to Chapter 6 of [Borkar, 2009].
@@
Using the log-likelihood trick [Williams, 1992] for the policy update, we have the gradient step
@@
Then following the characterization made by [Tessler et al., 2018] on the internally chain transitive invariant sets of the ODE Eq. 10, we can conclude the convergence for the two-timescale stochastic approximation processes. ∎
@@
See proof of D.1 in the Corollary. 30.2.2 of [Rockafellar, 1997]. For the proof of concavity of the perturbed C-MORL formulation $P(\xi)$, it suffices to show for every $\xi^{1},\xi^{2}\in\mathbb{R}^{n}$, and for any $\mu\in(0,1)$, the following equation holds:
@@
We refer the detailed proof of Eq. 12 in the Proposition 1 of [Paternain et al., 2019].
@@ E.1 Benchmark
To evaluate the performance of our method and baselines. we collect benchmarks from MO-Gymnasium Felten et al. [2023a] and SustainGym Yeh et al. [2024]. MO-Gymnasium is an open source Python library that includes more than 20 environments from diverse domains. SustainGym provides standardized benchmarks for sustainable energy systems, encompassing electric vehicles, data centers, electrical markets, power plants, and building energy systems. Among these environments, the building thermal control tasks involve large commercial buildings with multiple zones, aiming to regulate temperature while minimizing energy costs, making them suitable for developing multi-objective control tasks. Additionally, we incorporate the objective of demand response (reducing peak demand) and extend its multi-objective version. The details of all benchmarks are as follows:
@@
Fruit-Tree: A discrete MORL benchmark with a full binary tree of depth $d$ with randomly assigned vectorial reward $\mathbf{r}\in\mathbb{R}^{6}$ on the leaf nodes Yang et al. [2019]. These rewards are related to six objectives, showing the amounts of six different nutrition facts of the fruits on the tree:{Protein, Carbs, Fats, Vitamins, Minerals, Water}. The goal of the MORL agent is to find a path on the tree from the root to a leaf node that maximizes utility for a given preference. The reward of six objectives is defined as:
@@
This control task has two conflicting objectives: forward speed and control cost.
@@
MO-Ant-2d Y 2 $\mathcal{S}\subseteq\mathbb{R}^{27}$ $\mathcal{A}\subseteq\mathbb{R}^{8}$
MO-Ant-3d Y 3 $\mathcal{S}\subseteq\mathbb{R}^{27}$ $\mathcal{A}\subseteq\mathbb{R}^{8}$
MO-Humanoid-2d Y 2 $\mathcal{S}\subseteq\mathbb{R}^{376}$ $\mathcal{A}\subseteq\mathbb{R}^{17}$
@@ E.2 Training Details
We run all the experiments on a cloud server including CPU Intel Xeon Processor and GPU Tesla T4. In the Pareto initialization stage, we use PPO algorithm implemented by Kostrikov [2018]. The PPO parameters are reported in Table 6 and Table 7. For constrained optimization, we adopt C-MORL-IPO method. The hyperparameters of C-MORL-IPO include:
@@
  • log barrier coef: the tunable parameter on the log barrier.
@@ F.1 C-MORL-CPO
Recall in Section 5.3, we formulate C-MORL-CPO by solving the following optimization problem:
$\pi_{r+1,i}=\arg\max_{\pi\in\Pi_{\theta}}$ $G_{l}^{\pi}$ (18)
s.t. $G_{i}^{\pi}\geq\beta G_{i}^{\pi_{r}}\;i=1,\ldots,n,i\neq l$
$D\left(\pi,\pi_{r}\right)\leq\delta.$
However, solving this problem requires evaluation of the constraint functions to determine whether a proposed point $\pi$ is feasible. Therefore, follow [Achiam et al., 2017], we replace the objectives functions with surrogate functions, which are easy to estimate from samples collected on $\pi_{r}$. To be specific, we solve the following optimization problem to approximately solve Eq. 4:
@@
However, Eq. 19 is impractical to be solved directly especially when the policy is parameterized as a neural network with high-dimensional parameter spaces. We follow [Achiam et al., 2017] to implement an approximation to the update Eq. 19 that can be computed efficiently:
@@
In our implementation, we adopt C-MORL-CPO to solve MORL tasks with two objectives, i.e., constraint optimization problem that only involves single constraint. Therefore, we can directly compute the dual variables $\lambda^{*}$ and $\nu^{*}$ with the analytical solution as follows Achiam et al. [2017]:
@@ F.2 C-MORL-IPO
Recall in Section 5.3, we augment the objective function of the objective that is being optimized with logarithmic barrier functions for other constrained objectives. Note that the logarithmic barrier functions can be integrated with any other policy optimization methods, in this paper, we follow [Liu et al., 2020] to integrate it with PPO [Schulman et al., 2017] for training our policies. Therefore, the surrogate objective becomes:
@@ G.1 Parameter Study for Number of extension policies
G.2 Expected Utility results

### v2

@@ 6 Experiments
Next, to better understand the influence of hyperparameters and key components of C-MORL, we perform an in-depth analysis and conduct ablation studies focusing on these aspects.
Environments Metrics CAPQL Q-Pensieve PG-MORL GPI-LS MORL/D C-MORL
MO-Hopper-2d HV$(10^{5})$ 1.15$\pm$0.08 1.26$\pm$0.01 1.20$\pm$0.09 1.19$\pm$0.10 1.11$\pm$0.03 1.39$\pm$0.01
EU$(10^{2})$ 2.28$\pm$0.07 2.28$\pm$0.01 2.34$\pm$0.10 2.33$\pm$0.10 2.19$\pm$0.04 2.56$\pm$0.02
SP$(10^{2})$ 0.46$\pm$0.10 1.61$\pm$1.31 5.13$\pm$5.81 0.49$\pm$0.37 2.72$\pm$2.05 0.33$\pm$0.28
MO-Hopper-3d HV$(10^{7})$ 1.65$\pm$0.45 1.66$\pm$1.20 1.59$\pm$0.45 1.70$\pm$0.29 1.94$\pm$0.05 2.29$\pm$0.01
EU$(10^{2})$ 1.53$\pm$0.28 1.26$\pm$0.79 1.47$\pm$0.25 1.62$\pm$0.10 1.68$\pm$0.02 1.80$\pm$0.01
SP$(10^{2})$ 2.31$\pm$3.16 1.77$\pm$0.88 0.76$\pm$0.91 0.74$\pm$1.22 0.84$\pm$0.17 0.28$\pm$0.09
MO-Ant-2d HV$(10^{5})$ 1.11$\pm$0.69 2.55$\pm$0.54 0.35$\pm$0.08 3.10$\pm$0.25 1.03$\pm$0.26 3.13$\pm$0.20
EU$(10^{2})$ 2.16$\pm$0.94 3.14$\pm$0.49 0.81$\pm$0.23 4.28$\pm$0.19 2.22$\pm$0.49 4.29$\pm$0.19
SP$(10^{3})$ 0.18$\pm$0.07 3.63$\pm$2.71 2.20$\pm$3.48 3.61$\pm$2.13 2.90$\pm$2.61 1.67$\pm$0.85
MO-Ant-3d HV$(10^{7})$ 1.22$\pm$0.33 3.82$\pm$0.43 0.94$\pm$0.12 0.55$\pm$0.81 1.00$\pm$0.17 4.09$\pm$0.13
EU$(10^{2})$ 1.30$\pm$0.29 2.18$\pm$0.41 1.07$\pm$0.07 2.41$\pm$0.20 1.51$\pm$0.12 2.57$\pm$0.06
SP$(10^{3})$ 0.17$\pm$0.09 0.83$\pm$0.07 0.02$\pm$0.01 1.96$\pm$0.79 0.85$\pm$0.57 0.03$\pm$0.01
MO-Humanoid-2d HV$(10^{5})$ 3.30$\pm$0.05 0.90$\pm$0.62 2.62$\pm$0.32 1.98$\pm$0.02 2.81$\pm$0.07 3.43$\pm$0.06
EU$(10^{2})$ 4.75$\pm$0.04 1.89$\pm$0.45 4.06$\pm$0.32 3.67$\pm$0.02 4.32$\pm$0.06 4.78$\pm$0.05
SP$(10^{4})$ 0.00$\pm$0.00 1.08$\pm$1.39 0.13$\pm$0.17 0.00$\pm$0.00 0.00$\pm$0.00 0.18$\pm$0.27
Building-3d HV$(10^{12})$ 0.33$\pm$0.18 1.00$\pm$0.02 0.83$\pm$0.02 0.26$\pm$0.04 0.87$\pm$0.38 1.15$\pm$0.00
EU$(10^{4})$ 0.75$\pm$0.09 0.96$\pm$0.00 0.93$\pm$0.01 0.74$\pm$0.01 0.95$\pm$0.00 1.02$\pm$0.00
SP$(10^{5})$ 0.18$\pm$0.08 0.92$\pm$0.78 0.04$\pm$0.02 0.07$\pm$0.09 7.31$\pm$2.20 0.07$\pm$0.06
Building-9d HV$(10^{31})$ 4.29$\pm$0.73 7.28$\pm$0.57 T/O T/O T/O 7.93$\pm$0.07
EU$(10^{3})$ 3.31$\pm$0.06 3.46$\pm$0.03 T/O T/O T/O 3.50$\pm$0.00
SP$(10^{3})$ 4.34$\pm$3.72 1.04$\pm$0.38 T/O T/O T/O 2.79$\pm$0.40
Table 3: Parameter study of C-MORL for Policy Initialization on MO-Hopper-2d benchmark.
HV$(10^{5})$ EU$(10^{2})$ SP$(10^{2})$
M=3 (1.5M steps) 1.38$\pm$0.12 2.53$\pm$0.13 0.58$\pm$ 0.43
M=6 (1.5M steps) 1.39$\pm$0.04 2.55$\pm$0.04 0.16$\pm$0.11
M=11 (1.5M steps) 1.32$\pm$0.03 2.47$\pm$0.03 0.44$\pm$0.33
M=6 (3M steps) 1.45$\pm$0.05 2.63$\pm$0.05 0.22$\pm$0.10
Parameter study of C-MORL for Pareto Initialization. As mentioned in Section 5.1, during Pareto initialization stage, C-MORL aims to derive a few Pareto optimal solutions by training a set of initial policies. To study the impact of the number of initial policies $M$, we conduct experiments while keeping the total number of training steps fixed at $1.5\times 10^{6}$ steps (including $1\times 10^{6}$ steps for initialization stage) to ensure a fair comparison. Specifically, we uniformly sample $M=3,6,11$ preference vectors, as shown in Table 3. For example, with a sampling interval of $0.2$, the preference vectors are $\bm{\omega}=[0,1],[0.2,0.8],[0.4,0.6],[0.6,0.4],[0.8,0.2],[1,0]$. Intuitively, increasing $M$ can enhance diversity among Pareto-optimal solutions, which benefits the subsequent extension phase. However, since the number of total time steps is fixed, increasing $M$ reduces the training steps allocated to each initial policy. The trade-off is evident as the performance for $M=11$ is worse than for $M=6$, indicating that the increasing of $M$ does not always guarantee better performance. To further investigate, we increase the number of training steps to $3\times 10^{6}$ (with $2\times 10^{6}$ steps allocated to initialization stage) while keeping $M=6$. The results demonstrate that proportionally increasing the total training steps and the number of initial policies leads to improved performance, highlighting the importance of balancing training resources with policy diversity.
@@ Appendix A Policy Selection and C-MORL Algorithm
In this Section, we provide a brief proof for the convergence of iterates majorly inspired by (Tessler et al., 2018). For detailed proof, we refer to Chapter 6 of (Borkar, 2009).
@@
Using the log-likelihood trick (Williams, 1992) for the policy update, we have the gradient step
@@
Then following the characterization made by (Tessler et al., 2018) on the internally chain transitive invariant sets of the ODE Eq. 9, we can conclude the convergence for the two-timescale stochastic approximation processes. ∎
@@
In D.1, Slater’s condition requires that there exists a feasible policy $\pi$ such that all inequality constraints are strictly satisfied for problem (1), i.e.,
$G_{i}^{\pi}>d_{i},\quad\forall i=1,\ldots,n,\;i\neq l.$
See proof of D.1 in the Corollary. 30.2.2 of (Rockafellar, 1997). For the proof of concavity of the perturbed C-MORL formulation $P(\xi)$, it suffices to show for every $\xi^{1},\xi^{2}\in\mathbb{R}^{n}$, and for any $\mu\in(0,1)$, the following equation holds:
@@
We refer the detailed proof of Eq. 11 in the Proposition 1 of (Paternain et al., 2019).
@@ E.1 Benchmark
To evaluate the performance of our method and baselines. we collect benchmarks from MO-Gymnasium Felten et al. (2023a) and SustainGym Yeh et al. (2024). MO-Gymnasium is an open source Python library that includes more than 20 environments from diverse domains. SustainGym provides standardized benchmarks for sustainable energy systems, encompassing electric vehicles, data centers, electrical markets, power plants, and building energy systems. Among these environments, the building thermal control tasks involve large commercial buildings with multiple zones, aiming to regulate temperature while minimizing energy costs, making them suitable for developing multi-objective control tasks. Additionally, we incorporate the objective of demand response (reducing peak demand) and extend its multi-objective version. The details of all benchmarks are as follows:
@@
Fruit-Tree: A discrete MORL benchmark with a full binary tree of depth $d$ with randomly assigned vectorial reward $\mathbf{r}\in\mathbb{R}^{6}$ on the leaf nodes Yang et al. (2019). These rewards are related to six objectives, showing the amounts of six different nutrition facts of the fruits on the tree:{Protein, Carbs, Fats, Vitamins, Minerals, Water}. The goal of the MORL agent is to find a path on the tree from the root to a leaf node that maximizes utility for a given preference. The reward of six objectives is defined as:
@@
This control task has two conflicting objectives: forward speed and control cost. MO-Humanoid was chosen because it has one of the most complex state space among all Mujoco environments, with 348 states.
@@
MO-Ant-2d Y 2 $\mathcal{S}\subseteq\mathbb{R}^{105}$ $\mathcal{A}\subseteq\mathbb{R}^{8}$
MO-Ant-3d Y 3 $\mathcal{S}\subseteq\mathbb{R}^{105}$ $\mathcal{A}\subseteq\mathbb{R}^{8}$
MO-Humanoid-2d Y 2 $\mathcal{S}\subseteq\mathbb{R}^{348}$ $\mathcal{A}\subseteq\mathbb{R}^{17}$
@@ E.2 Training Details
We run all the experiments on a cloud server including CPU Intel Xeon Processor and GPU Tesla T4. In the Pareto initialization stage, we use PPO algorithm implemented by Kostrikov (2018). The PPO parameters are reported in Table 7 and Table 8. For constrained optimization, we adopt C-MORL-IPO method. For the baseline implementations, Q-Pensieve (Hung et al., 2022) and PG-MORL (Xu et al., 2020) are reproduced using the official code provided by their respective papers, ensuring consistency with the original experiments. For Envelope (Yang et al., 2019), CAPQL (Lu et al., 2022), GPI-LS (Alegre et al., 2023), and MORL/D (Felten et al., 2024), we utilize the implementations available in the MORL-baselines library (Felten et al., 2023a), adapting them as necessary to align with our experimental setup. The hyperparameters of C-MORL-IPO include:
@@
  • Log barrier coef $t$: the tunable parameter on the log barrier.
@@
In policy initialization stage, the preference vectors for training initial policies are uniformly sampled from the preference space $\Omega$. For example, in the case of MO-Ant-2d, we set the sampling interval to $0.2$, with the minimum and maximum values for each dimension being 00 and $1$, respectively. This results in the following preference vectors: $[0,1],[0.2,0.8],[0.4,0.6],[0.6,0.4],[0.8,0.2]$, and $[1,0]$, for a total of 6 preference vectors.
@@ F.1 C-MORL-CPO
C-MORL-CPO Rather than applying sampling-based approaches (Duan et al., 2016) for finding policy updates in the relaxed formulation, empirically we can follow the adapted trust region method (Schulman et al., 2015) for CPO updates (Achiam et al., 2017):
However, solving this problem requires evaluation of the constraint functions to determine whether a proposed point $\pi$ is feasible. Therefore, follow (Achiam et al., 2017), we replace the objectives functions with surrogate functions, which are easy to estimate from samples collected on $\pi_{r}$. To be specific, we solve the following optimization problem to approximately solve Eq. 17:
@@
However, Eq. 18 is impractical to be solved directly especially when the policy is parameterized as a neural network with high-dimensional parameter spaces. We follow (Achiam et al., 2017) to implement an approximation to the update Eq. 18 that can be computed efficiently:
@@
In our implementation, we adopt C-MORL-CPO to solve MORL tasks with two objectives, i.e., constrained optimization problem that only involves single constraint. Therefore, we can directly compute the dual variables $\lambda^{*}$ and $\nu^{*}$ with the analytical solution as follows Achiam et al. (2017):
@@ F.2 C-MORL-IPO
Recall in Section 5.3, we augment the objective function of the objective that is being optimized with logarithmic barrier functions for other constrained objectives. Note that the logarithmic barrier functions can be integrated with any other policy optimization methods, in this paper, we follow (Liu et al., 2020) to integrate it with PPO (Schulman et al., 2017) for training our policies. Therefore, the surrogate objective becomes:
@@ G.1 Parameter Study for Number of extension policies
G.2 Ablation Study on Buffer Size.
In order to better understand how Buffer size can influence the performance of C-MORL, we provide the comparison of the best baseline and C-MORL with various buffer sizes, as shown in Table 12 and Table 13. These studies show that even with reduced buffer sizes, C-MORL maintains highly competitive performance. Compared to PG-MORL, our C-MORL achieves superior results while requiring significantly fewer policies. Another notable observation is that, in some benchmarks, the results remain consistent regardless of the buffer size. This consistency indicates that the number of Pareto optimal policies in these cases does not exceed the buffer size.
Table 12: Evaluation of HV, EU, and SP in discrete MORL tasks under varying buffer sizes, alongside a comparison with the best-performing baseline.
Environments Metrics Best Baseline B=20 B=50 B=100 B=200
Minecart HV$(10^{2})$ 6.05$\pm$0.37(GPI-LS) 6.22$\pm$0.70 6.57$\pm$0.92 6.63$\pm$0.89 6.77$\pm$0.88
EU$(10^{-1})$ 2.29$\pm$0.32(GPI-LS) 1.88$\pm$0.59 2.05$\pm$0.67 2.09$\pm$0.65 2.12$\pm$0.66
SP$(10^{-1})$ 0.10$\pm$0.00(GPI-LS) 0.64$\pm$0.17 0.19$\pm$0.02 0.09$\pm$0.03 0.05$\pm$0.02
MO-Lunar-Lander HV$(10^{9})$ 1.06$\pm$0.16(GPI-LS) 1.01$\pm$0.07 1.08$\pm$0.04 1.12$\pm$0.03 1.12$\pm$0.03
EU$(10^{1})$ 1.81$\pm$0.34(GPI-LS) 1.84$\pm$0.31 2.21$\pm$0.23 2.35$\pm$0.18 2.35$\pm$0.18
SP$(10^{3})$ 0.13$\pm$0.01(GPI-LS) 1.14$\pm$2.14 1.65$\pm$1.83 1.04$\pm$1.24 1.04$\pm$1.24
Fruit-Tree HV$(10^{4})$ 3.66$\pm$0.23(Envelope) 2.34$\pm$0.29 2.86$\pm$0.19 3.17$\pm$0.20 3.67$\pm$0.14
EU 6.15±0.00(Envelope/GPI-LS) 6.14$\pm$0.13 6.38$\pm$0.10 6.46$\pm$0.08 6.53$\pm$0.03
SP 0.53±0.02(GPI-LS) 2.62$\pm$0.49 1.65$\pm$1.83 0.81$\pm$0.12 0.04±0.00
Table 13: Evaluation of HV, EU, and SP in continuous MORL tasks under varying buffer sizes, alongside a comparison with the best-performing baseline.
Environments Metrics Best Baseline B=20 B=50 B=100 B=200
MO-Hopper-2d HV$(10^{5})$ 1.26$\pm$0.01(Q-Pensieve) 1.39$\pm$0.01 1.39$\pm$0.01 1.39$\pm$0.01 -
EU$(10^{2})$ 2.34$\pm$0.10(PG-MORL) 2.55$\pm$0.01 2.56$\pm$0.02 2.56$\pm$0.02 -
SP$(10^{2})$ 0.46$\pm$0.10(CAPQL) 2.68$\pm$1.66 0.57$\pm$0.29 0.33$\pm$0.28 -
MO-Hopper-3d HV$(10^{7})$ 1.70$\pm$0.29(GPI-LS) 2.03$\pm$0.15 2.20$\pm$0.04 2.26$\pm$0.02 2.29$\pm$0.01
EU$(10^{2})$ 1.62$\pm$0.10(GPI-LS) 1.72$\pm$0.08 1.78$\pm$0.02 1.80$\pm$0.01 1.80$\pm$0.01
SP$(10^{2})$ 0.74$\pm$1.22(GPI-LS) 7.61$\pm$2.24 2.74$\pm$1.91 0.86$\pm$0.27 0.28$\pm$0.09
MO-Ant-2d HV$(10^{5})$ 3.10$\pm$0.25(GPI-LS) 3.08$\pm$0.21 3.13$\pm$0.20 3.13$\pm$0.20 -
EU$(10^{2})$ 4.28$\pm$0.19(GPI-LS) 4.27$\pm$0.19 4.29$\pm$0.19 4.29$\pm$0.19 -
SP$(10^{3})$ 0.18$\pm$0.07(CAPQL) 3.66$\pm$1.29 1.67$\pm$0.85 1.67$\pm$0.85 -
MO-Ant-3d HV$(10^{7})$ 3.82$\pm$0.43(Q-Pensieve) 3.52$\pm$0.16 3.83$\pm$0.17 4.00$\pm$0.12 4.09$\pm$0.13
EU$(10^{2})$ 2.41$\pm$0.20(GPI-LS) 2.48$\pm$0.09 2.55$\pm$0.08 2.57$\pm$0.07 2.57$\pm$0.06
SP$(10^{3})$ 0.02$\pm$0.01(PG-MORL) 1.56$\pm$0.67 0.30$\pm$0.19 0.08$\pm$0.05 0.03$\pm$0.01
MO-Humanoid-2d HV$(10^{5})$ 3.30$\pm$0.05(CAPQL) 3.43$\pm$0.06 3.43$\pm$0.06 3.43$\pm$0.06 -
EU$(10^{2})$ 4.75$\pm$0.04(CAPQL) 4.78$\pm$0.05 4.78$\pm$0.05 4.78$\pm$0.05 -
SP$(10^{4})$ 0$\pm$0.00(CAPQL/GPI-LS) 0.18$\pm$0.27 0.18$\pm$0.27 0.18$\pm$0.27 -
Building-3d HV$(10^{12})$ 1.00$\pm$0.02(Q-Pensieve) 1.14$\pm$0.00 1.14$\pm$0.00 1.15$\pm$0.00 1.15$\pm$0.00
EU$(10^{4})$ 0.96$\pm$0.00(Q-Pensieve) 1.02$\pm$0.00 1.02$\pm$0.00 1.02$\pm$0.00 1.02$\pm$0.00
SP$(10^{4})$ 0.37$\pm$0.22(PG-MORL) 1.04$\pm$0.16 1.98$\pm$0.38 0.69$\pm$0.62 0.69$\pm$0.62
Building-9d HV$(10^{31})$ 7.28$\pm$0.57(Q-Pensieve) 7.64$\pm$0.17 7.93$\pm$0.07 7.93$\pm$0.07 7.93$\pm$0.07
EU$(10^{3})$ 3.46$\pm$0.03(Q-Pensieve) 3.50$\pm$0.00 3.52$\pm$0.00 3.52$\pm$0.00 3.52$\pm$0.00
SP$(10^{4})$ 0.10$\pm$0.04(Q-Pensieve) 1.16$\pm$0.19 0.28$\pm$0.03 0.28$\pm$0.04 0.28$\pm$0.04
G.3 Expected Utility results